# Generalized Finite-time Optimal Control Framework in Stochastic Thermodynamics

### v3

@@ 5.2 Stochastic particle in a harmonic trap and
A stochastic particle in a harmonic trap potential models an important equivalence class that describes a large class of physical systems, and hence has been the testbed of novel theoretical developments in ST. The stochastic particle satisfies the Boltzmann distribution, $P\propto e^{-k(x-m)^{2}}$, with the trap stiffness $k$ and center $m$. In this case, these two relevant control parameters fully capture the statistical properties of the system. We consider a finite-time optimal control problem for the time-dependent change of the trap stiffness $k(t)$ and center $m(t)$ [10, 24]. The inverse of the trap stiffness quantifies the covariance $(C)$ of the Gaussian distribution for the particle position, $C\propto 1/k$ [10]. Hence, by definition, the finite-time optimal control formulation for changing the centre and stiffness of the harmonic trap is equivalent to the optimal transport theory to change the mean and covariance of the Gaussian distribution for the position of the particles [49] or the application of thermodynamic geometry to the harmonic trap [24]. However, the results obtained from different methods do not match [10, 24, 49, 58]. We resolve these discrepancies using the GFTOC framework.
@@ 6 Conclusion and Outlook
Our framework opens up a plethora of practical applications in biology, chemistry, population dynamics, and nanoscale / mesoscale devices, where stochastic thermodynamics has been an experimentally tested theoretical paradigm [4]. Our framework opens the door for the experimental design, optimization, and control of such systems. The experimental verification of our framework awaits exploration [37]. The GFTOC framework is rather trivially extended to quantum systems, as the development of optimal control and geometric speed limits for quantum systems is attributed to the underlying geometric structure and not to the quantum nature [91, 92, 93]. More broadly, the GFTOC methodology developed here in stochastic thermodynamics applies to other interdisciplinary domains that rely on Riemannian geometry as their mathematical foundation, including information geometry [38], in which the role of ‘kinks’ warrants further investigation.

### v4

@@ 5.2 Stochastic particle in a harmonic trap and
A stochastic particle in a harmonic trap potential models an important equivalence class that describes a large class of physical systems, and hence has been the testbed of novel theoretical developments in ST. The stochastic particle satisfies the Boltzmann distribution, $P\propto e^{-k(x-m)^{2}}$, with the trap stiffness $k$ and center $m$. In this case, these two relevant control parameters fully capture the statistical properties of the system. We consider a finite-time optimal control problem for the time-dependent change of the trap stiffness $k(t)$ and center $m(t)$ [10, 24]. The inverse of the trap stiffness quantifies the covariance $(C)$ of the Gaussian distribution for the particle position, $C\propto 1/k$ [10]. Hence, by definition, the finite-time optimal control formulation for changing the center and stiffness of the harmonic trap is equivalent to the optimal transport theory to change the mean and covariance of the Gaussian distribution for the position of the particles [49] or the application of thermodynamic geometry to the harmonic trap [24]. However, the results obtained from different methods do not match [10, 24, 49, 58]. We resolve these discrepancies using the GFTOC framework.
@@ 6 Conclusion and Outlook
Our framework opens up a plethora of practical applications in biology, chemistry, population dynamics, and nanoscale / mesoscale devices, where stochastic thermodynamics has been an experimentally tested theoretical paradigm [4]. Our framework opens the door for the experimental design, optimization, and control of such systems. The experimental verification of our framework awaits exploration [37]. The GFTOC framework is rather trivially extended to quantum systems, as the development of optimal control and speed limits for quantum systems is attributed to the underlying geometric structure and not to the quantum nature [91, 92, 93]. More broadly, the GFTOC methodology developed here in stochastic thermodynamics applies to other interdisciplinary domains that rely on Riemannian geometry as their mathematical foundation, including information geometry [38], in which the role of ‘kinks’ warrants further investigation.